40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Mon May 24 15:47:30 2021, Radhika, DailyProgress, Cryo vacuum chamber, Weekend cryo chamber cooling cooling.png
    Reply  Mon May 24 16:49:17 2021, rana, DailyProgress, Cryo vacuum chamber, Weekend cryo chamber cooling PPP.jpg
       Reply  Tue May 25 13:19:41 2021, Radhika, DailyProgress, Cryo vacuum chamber, Weekend cryo chamber cooling 
    Reply  Tue May 25 16:38:17 2021, Stephen, DailyProgress, Cryo vacuum chamber, Weekend cryo chamber cooling --> started warmup cooling_full_20210525.png
       Reply  Thu May 27 16:48:35 2021, Radhika, DailyProgress, Cryo vacuum chamber, Weekend cryo chamber cooling --> started warmup 7x
          Reply  Mon May 31 11:38:13 2021, Radhika, DailyProgress, Cryo vacuum chamber, Weekend cryo chamber cooling 14x
             Reply  Wed Jun 2 20:06:07 2021, rana, DailyProgress, Cryo vacuum chamber, Weekend cryo chamber cooling 
Message ID: 2580     Entry time: Mon May 31 11:38:13 2021     In reply to: 2579     Reply to this: 2581
Author: Radhika 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Weekend cryo chamber cooling 

[Stephen, Radhika]

After Thursday's work, we resumed on Friday and lifted up the coldplate to access the collar and baseplate. Stephen added metallized mylar wraps to the peek cylindrical spacers [pic 9, 10]. He took out the cylindrical collar and baseplate (previously there to minimize contact between the colplate and chamber bottom) and replaced them with a foil collar [pic 13, note this is before pushing down the foil over the PEEK spacers]. We re-inserted the coldplate.

In order to wrap/insulate the copper braid, we inserted 2 sheets of metallized mylar into the vacuum tube to surround the braid [pic 11, 14]. We cut the mylar wrapping so that it did not short to the outer shield. We also confirmed that the copper braid is not shorting to the inner shield hole (there was clearance for the shield to be lifted before hitting the braid). The mylar extends to the coldhead, which we then wrapped with aluminum foil, making sure not to short to the walls of the tube [pic 12]. 

We switched the foil wrapping from the inside to the outside of the outer shield, so that any radiative transfer from the inner shield to the outer shield would be absorbed as much as possible (not reflected back). We placed both shields back and bolted down the copper braid loop and workpiece. We re-attached the RTDs, then placed in both shield lids (without bolting them down) and stopped for the day.

Today I checked on the chamber setup and closed up. I started the vacuum pump and it made abnormally loud noises, indicating something was off. The percentage sign continued to flash, indicating that the pump was not reaching 80% speed. I performed 2-3 power cycles, which did not solve the issue. We will pick up to debug the issue early this week.

[update 02 June 2021]

Pumpdown and cooldown were successfully started this morning - Radhika retightened the green leak valve and pumps started just fine. We will check trends on Friday and likely will allow cooldown to proceed over the weekend.

We are both working on adding all of our photos to the photo dump at the ligo.wbridge QIL Cryostat Photo Album.

Attachment 1: IMG_1681.jpeg  2.062 MB  Uploaded Mon May 31 20:55:14 2021  | Hide | Hide all
IMG_1681.jpeg
Attachment 2: IMG_1684.jpeg  2.147 MB  Uploaded Mon May 31 20:55:30 2021  | Hide | Hide all
IMG_1684.jpeg
Attachment 3: IMG_1686.jpeg  2.103 MB  Uploaded Mon May 31 20:55:45 2021  | Hide | Hide all
IMG_1686.jpeg
Attachment 4: IMG_1687.jpeg  1.744 MB  Uploaded Mon May 31 20:56:04 2021  | Hide | Hide all
IMG_1687.jpeg
Attachment 5: IMG_1688.jpeg  1.383 MB  Uploaded Mon May 31 20:56:29 2021  | Hide | Hide all
IMG_1688.jpeg
Attachment 6: IMG_1690.jpeg  2.347 MB  Uploaded Mon May 31 21:00:31 2021  | Hide | Hide all
IMG_1690.jpeg
Attachment 7: IMG_1693.jpeg  2.037 MB  Uploaded Mon May 31 21:00:46 2021  | Hide | Hide all
IMG_1693.jpeg
Attachment 8: IMG_1695.jpeg  1.802 MB  Uploaded Mon May 31 21:02:11 2021  | Hide | Hide all
IMG_1695.jpeg
Attachment 9: IMG_8796.JPG  2.617 MB  Uploaded Wed Jun 2 19:17:57 2021  | Hide | Hide all
IMG_8796.JPG
Attachment 10: IMG_8805.JPG  2.352 MB  Uploaded Wed Jun 2 19:19:30 2021  | Hide | Hide all
IMG_8805.JPG
Attachment 11: IMG_8811.JPG  2.703 MB  Uploaded Wed Jun 2 19:20:10 2021  | Hide | Hide all
IMG_8811.JPG
Attachment 12: IMG_8815.JPG  3.362 MB  Uploaded Wed Jun 2 19:20:40 2021  | Hide | Hide all
IMG_8815.JPG
Attachment 13: IMG_8803.JPG  3.333 MB  Uploaded Wed Jun 2 19:22:35 2021  | Hide | Hide all
IMG_8803.JPG
Attachment 14: IMG_8810.JPG  2.896 MB  Uploaded Wed Jun 2 19:24:43 2021  | Hide | Hide all
IMG_8810.JPG
ELOG V3.1.3-